OpenStudy (anonymous):

Explain why the natural log function y=lnx used much more frequently in calculus than the other log functions y=log_a (x)?

OpenStudy (anonymous):

I was thinking because when you inverse the y=lnx, to e^y=x you get the special case, that the derivative of e^x=e^x (equals itself)

OpenStudy (anonymous):

is there anything else to it ?

OpenStudy (gorv):

because standard values are defined for log base 10 so other values are not used